OpenStudy (anonymous):

Can't get this question correct: (2^3 + ^2)(2^3-^2)

OpenStudy (anonymous):

\[(2\sqrt{3} + \sqrt{2}) (2\sqrt{3}-\sqrt{2})\]

OpenStudy (anonymous):

@jim_thompson5910 can you help me

hartnn (hartnn):

u know the formula : \((a+b)(a-b)=a^2-b^2\)

OpenStudy (anonymous):

yes but i don't understand why answer 10

hartnn (hartnn):

did u try to apply that formula here ? what did u get ?

hartnn (hartnn):

a=\(2\sqrt3\\b=\sqrt2\)

OpenStudy (anonymous):

I got that but do you multiply the ^3

OpenStudy (anonymous):

i got it

hartnn (hartnn):

what did u get ?

OpenStudy (anonymous):

\[4\sqrt{9}-2\sqrt{6}+2\sqrt{6}-\sqrt{4}=12-2=10\]

hartnn (hartnn):

u are correct! good work :)

OpenStudy (anonymous):

you really helped me visuALIZE IT THANKS

hartnn (hartnn):

welcome :)